OpenStudy (needshelpdfdf):

Tiles 2x2 − 32 = 0 4x2 − 100 = 0 x2 − 55 = 9 x2 − 140 = -19 2x2 − 18 = 0 Pairs {-8, 8} arrowBoth {-4, 4} arrowBoth {-5, 5} arrowBoth {-11, 11} arrowBoth

OpenStudy (alivejeremy):

2x2 − 32 = 0

OpenStudy (alivejeremy):

Is {-4, 4}

OpenStudy (alivejeremy):

4x2 − 100 = 0

OpenStudy (alivejeremy):

is {-5, 5}

563blackghost (563blackghost):

Lets do the first one first.... \(\huge\bf{2x^{2}-32=0}\) We need the variable \(\large{2x^{2}}\) to itself so we would add 32 to each side.... \(\huge{2x^{2}=32}\) Now we need to divide by 2 from each side... \(\huge{{x^{2}=16}}\) In which we find the square root... \(\huge{\sqrt{16}=x}\)

OpenStudy (alivejeremy):

x2 − 55 = 9

OpenStudy (alivejeremy):

is {-8, 8}

563blackghost (563blackghost):

@alivejeremy you can not be giving the answers away, you must help the owlet so that they may understand the question and how to solve. This is in violation of the code of conduct. Please refrain from this in the future.

OpenStudy (needshelpdfdf):

black stop being a wingspan head

OpenStudy (alivejeremy):

x2 − 140 = -19

OpenStudy (alivejeremy):

is

563blackghost (563blackghost):

Excuse me?

OpenStudy (ineedhelpduh):

ho stfu

OpenStudy (alivejeremy):

{-11, 11}

OpenStudy (alivejeremy):

2x2 − 18 = 0

OpenStudy (alivejeremy):

is

OpenStudy (alivejeremy):

{-3, 3}

OpenStudy (needshelpdfdf):

thank you alive

OpenStudy (alivejeremy):

Np.

563blackghost (563blackghost):

If you are given the answer you'll never learn on how to solve them.

OpenStudy (needshelpdfdf):

ok cool you can leave black